### v1

@@ 3.4 Coupling between different solution spaces
This section details the implementation issue of mixing solutions from different solution spaces $V_{h}$ and $W_{h}$. In a previous work [34], the coupling is done through the projection of discontinuous $\zeta$ into the continuous space $W_{h}$, defined by finding $\eta\in W_{h}$ satisfying:
@@ 3.5 Summary of DG and CG discretization
DG_Timestep
@@
9:Accumulate the quadrature $\mathbf{L}_{k}^{e^{+}}$ using $\widehat{\mathbf{F}}\cdot\mathbf{n}\psi_{k}$
[…]
@@
CG_Momentum_Timestepsseach element $e$
[…]
8:Compute elemental $\alpha,\beta_{\text{barotropic}}$ from $\eta^{s}_{i},\eta^{s+1}_{i},P^{s}(\mathbf{x}_{i}),P^{s+1}(\mathbf{x}_{i}),\;i=1,2,3$
[…]
12:Compute nodal $\alpha,\beta_{\text{fric}}$ using $H^{s}_{i},H^{s+1}_{i},u_{x}^{s},u_{x}^{s+1},u_{y}^{s},u_{y}^{s+1}$
13:Compute nodal $\alpha,\beta_{\text{coriolis}}$ using $u_{x}^{s},u_{x}^{s+1},u_{y}^{s},u_{y}^{s+1}$
[…]
16:Compute $\gamma_{\text{fric}},\gamma_{\text{coriolis}}$
[…]
@@ 4.2 DG approach
Wetting and drying in DG-SWEM is considerably simpler as we only have to deal with elements and not nodes [23]. It is constructed such that elemental mass is always conserved, and momentum whenever possible. For each element we apply the PositiveDepth() procedure which performs the following:
@@ 4.3 Combined approach
PositiveDepth each element $e$
@@
[…]
@@
[…]
@@ 5 Incorporating Rainfall Data
During tropical cyclones, rainfall patterns in the vicinity of the storm exhibit defined structures that can be exploited by so-called parametric rainfall models. The basic idea behind this approach is to construct rainfall fields using simple analytic expressions based on a small number of (predicted or observed) storm parameters that are made publicly available by the National Hurricane Center (NHC) pre and post storm events. This type of approach has been implemented in the code based on the so-called R-CLIPER (Rainfall CLImatology and PERsistence) and the Interagency Performance Evaluation Task Force Rainfall Analysis (IPET) model, which computes rainfall rates $R(r)$ at any point in the grid, where $r$ is the distance from that point to the storm center. As an example, the R-CLIPER datapath is illustrated in Figure 3; see [35, 36] for more details on general parametric schemes. Note that the primary inputs of the model are the reported storm centers and maximum wind velocities, which are generally provided at 6-hour intervals by the NHC. Within DG-SWEM, linear interpolation is used to obtain the storm parameters at the model time step, and the rainfall rate at each finite element mesh node is computed based on its radial distance from the storm center.
@@
Alternatively, the program should also accept observed or reanalysis rainfall data in hindcasting scenarios in the GRIB2 format [37] which is read during each timestep. As this type of data is given with spatially varying distributions over the ocean and land, this data is readily interpolated onto finite element meshes. This option has been implemented in DG-SWEM and will be incorporated into ADCIRC in the future.
6 Integration into the ADCIRC codebase
This section gives a high-level overview of how the code is organized and designed. A detailed overview of ADCIRC architecture itself is shown in . As mentioned, the DG code was originally written as a separate software DG-SWEM; a robust integration of the relevant parts into the ADCIRC code is a major part of this work. Several key points were considered in this process:
  1. 1.
    ADCIRC must remain fully backward compatible. Unless the DG solver is chosen, the behavior of the code is not changed in any way.
  2. 2.
    Coupling between the ADCIRC code and DG code is minimized.
  3. 3.
    Redundancy of variables representing the same quantities, e.g. $h_{b}$, is minimized.
The first approach that was considered was to loosely couple the two programs by wrapping the calls to each solver in a high-level scripting language. A good choice was Python through its support of interfacing to Fortran using F2PY [^1] https://numpy.org/doc/stable/f2py/index.html. Using this approach, data arrays and subroutines can be accessed through NumPy and communicated back and forth between DG and CG. This approach was used in [38] to couple ADCIRC with other hydrodynamic software packages such as GSSHA and AdH. The main advantage of this approach is modularity as it treats each solver as a standalone module, and no modifications to the source code are required if we use the automatic wrapper generation of F2PY. However, there are several issues with this approach, particularly for this work. Because ADCIRC and DG-SWEM have many overlapping data variables such as bathymetric data, input options, mesh connectivity, such loose coupling would lead to wasteful space usage and redundant reading of input files. Attempting to eliminate this redundancy by sharing the overlapping data variables would then introduce a deeper coupling. Furthermore, using a wrapper presents a new interface which will require work to be adopted to existing workflows.
The current approach embeds the DG solver into a separate module within ADCIRC. A basic module dependency diagram is shown in Figure 4. First, we manually identify the common data arrays and variables between ADCIRC and DG-SWEM (with help from compiler error messages), then redirect the imports of those in DG to the appropriate modules in ADCIRC, such as global.F and mesh.F. This keeps the size of the DG data module small and thus addresses aim (3). Next, coupling from ADCIRC to DG is minimized since the only dependence is through calling the DG solver in the main timestepping loop. The DG solver essentially maintains the same interface as the original GWCE solver. This maintains modularity of the specific continuity solver and its separate from the main routines. Along with the DG module depending only on the basic low-level ADCIRC modules, this satisfies aim (2). Consequently, the first aim of backward compatibility is trivial since there is only one point of dependency on DG in ADCIRC. The option to use the DG solver is implemented as a new mode number in the existing ADCIRC configuration format.
Figure 4: Module dependency diagram of the DG solver in the context of ADCIRC. Arrow indicates that the pointer depends on the pointed. For clarity, we omit several modules from ADCIRC and show only the commonly used GLOBAL and MESH modules. In the time stepping loop, either the DG or GWCE solver is chosen based on an input parameter.
7 Numerical Experiments and Evaluation
@@
In this section, we thoroughly evaluate the developed solver by performing extensive numerical experiments that are designed to validate the solver. First, we consider simplified test cases where analytic solutions exist in simple rectangular and annular geometries. This is followed by the classical Shinnecock Inlet test case, a modified lake-at-rest [39] test case, and a relatively new compound flood modeling benchmark of the Neches River. Last, we perform large-scale validation test cases for Hurricanes Ike (2008) and Harvey (2017) where we simulate the entire western North Atlantic Ocean, Caribbean Sea, and Gulf of Mexico.
@@ 7.1 Lynch and Gray test case
To verify the numerical accuracy of the scheme, we investigate the convergence of the error as we refine the mesh spacing $h$. This test case was first described in [40] which also presented an analytical steady-state solution to be used in the verification. This is for the linearized version of the SWE, obtained by neglecting the advection terms and linearizing the friction term:
@@ 7.5 Neches River
Figure 13: Computational domain of the Neches river case. Observation gauges are marked as red circles and boundaries are highlighted in yellow [6].
@@
The next test case is based on an approach to model compound flooding through river-coastal ocean interaction. In this case, we consider the Lower Neches River watershed near the Texas-Louisiana border as shown in Figure 13. During Hurricane Harvey (2017), this watershed was subjected to extreme rainfall and large portions of its surrounding areas were inundated during and after the event, see e.g., [3] for details on the flooding during this event. The model domain has been extracted from the ADCIRC mesh used in the Hurricane Harvey case in Section 7.6 [41], and contains 122,839 elements and 62,075 nodes and is shown in Figure 13. Hence, the highly detailed unstructured mesh, bathymetry and topography, as well as parameters such as Manning’s $n$ are preserved. This event was also studied using ADCIRC and compared to HEC-RAS in [6] and the present study uses the same input data used in Loveland et al.. We only present key features of this model and refer readers to the original publication and [25] where we extended the DG-SWEM model for greater detail.
As in [6], river flow data was extracted from a validated HEC-RAS model and a USGS gauge and implemented as flux boundary condition at the upstream. At the downstream end of the river which terminates into Sabine Lake, a time-varying elevation boundary condition was applied based on the same HEC-RAS model and the closest NOAA gauge. The locations of these inputs and the gauges used for validation are shown in Figure 13.
Comparisons of outputs from CG, DG-CG, as well as observed data from August 20 to September 1, 2017 are shown in Figure 14. We observe underprediction from the CG solver in all cases except at the Rainbow Bridge. We also observe that in those cases, the output from DG-CG is higher and closer to peak observations. At Rainbow Bridge, both solvers match closely (while slightly overpredicting) up to the end of observed data. That the DG scheme is able to better handle flows with high advection than ADCIRC has also been demonstrated in other studies [24, 25]. The mismatch between wet/dry elevation values at the beginning of each solver can be attributed to differences in wetting-and-drying criteria as described in Section 4.
@@ 7.6 Hurricane Ike (2008) and Hurricane Harvey (2017)
In Figure 15, the computational mesh used is shown. This mesh consists of 2 million nodes and contains a recently updated version of the detailed floodplain, with a smallest element size of 120 meters [42]. Manning’s $n$ friction parameter is variable throughout the domain and was determined from observed datas [42]. As shown in Figure 15, the mesh is highly unstructured with increased refinement in the Texas-Louisiana coastal region. For Hurricane Ike, we use a proprietary wind data in the OWI format from OceanWeather Inc., see Section 4.1 in [41] for further details on the wind data. For Hurricane Harvey, we use a more refined version of the mesh which contains around 8 million nodes; this is needed to cover the floodplain region where Harvey primarily affected. Wind data in this case is obtained from the NHC Hurricane Database (HURDAT) in the Best Track format. Lastly, for both hurricanes, we obtain tides using OceanMesh2D [43], which in turn utilizes the TPXO9 tidal model [44] to define a periodic forcing based on the Q1, O1, P1, K1, N2, M2, S2, and K2 tidal constituents. Additionally, we apply a elevation boundary condition corresponding to these tides along the 60 meridian, i.e., the eastern portion of the mesh shown in Figure 15.
@@
Water elevation comparison for the Harvey case is shown in in Figure 17. We can clearly observe the effect of rain in the DG-CG case which raises the water level following the peak surge. This is most apparent at Aransas Wildlife Refuge (8774230) and Sargent (8772985). In the former, the secondary peak is delayed in DG-CG but matches the actual peak level better than CG without rain. In the latter, both solvers underpredict initial water levels, but DG-CG eventually matches the peak level. On the other hand, the parametric rain approach can also lead to overprediction as seen in the Packery Channel station. We also compare the simulated maximum water levels with the observed high water marks (HWM) [45] located in the region around Harvey’s landfall (Figure 18). Note that data counts differ between CG and DG-CG because we only count nodes that are wet. We see a marked improvement in the DG-CG case, indicating the impact of precipitation in compound flooding. A limitation in DG-CG wetting and drying for this case is that it was necessary to turn off momentum advection terms to prevent instability. Although these terms do not make a significant impact in this case, we do aim to make the DG-CG wetting and drying scheme more robust so as to handle highly refined meshes such as this one.
@@ 8.1.1 Optimization
As mentioned, the higher degrees of freedom in the DG method unavoidably requires more time to compute than the CG scheme. On the other hand, the higher degree of data parallelism from having most computations be element-local makes it attractive to use vector instructions. The general idea and motivation is outlined in [46] which involves significant loop reordering and blocking. To keep the code maintainable as part of ADCIRC, we opt to use a simpler approach and try to rely on automatic vectorization and basic optimizations as much as possible [47]. Wherever possible, loops with independent iterations and large trip count (e.g. number of elements) are prepended with the OpenMP !\$omp simd pragma, and inner loops are unrolled. The compiled code and assembly is checked using Intel Advisor to verify if the vectorization took place. Vectorizing the slope limiter requires a little more effort, as the automatic vectorizations were only partial. We manually performed explicit blocking on the number of elements and created vector versions of the local arrays private to each element. Without significant restructuring, the interior edge integration routine is not very suitable to vectorization since multiple edges can modify the same element. Nevertheless, we still observe significant improvement from inlining the numerical flux function by moving it into the same module.

### v2

@@ 3.4 Coupling between different solution spaces
This section details the implementation issue of mixing solutions from different solution spaces $V_{h}$ and $W_{h}$. In a previous work [Dawson2006-ve], the coupling is done through the projection of discontinuous $\zeta$ into the continuous space $W_{h}$, defined by finding $\eta\in W_{h}$ satisfying:
@@ 3.5 Summary of DG and CG discretization
1: procedure DG_Timestep
[…]
3:   for each edge $l$ do
4:    $e^{+}$ := right element of $l$
5:    $e^{-}$ := left element of $l$
6:    for each quadrature point $i$ do
@@
11:      Compute $\widehat{\mathbf{F}}:=\widehat{\mathbf{F}}(U,V,\eta^{+},\eta^{-},h_{b})$
12:      for each mode $k=1,2,3$ do
13:       Accumulate the quadrature $\mathbf{L}_{k}^{e^{+}}$ using $\widehat{\mathbf{F}}\cdot\mathbf{n}\psi_{k}$
14:       Accumulate the quadrature $\mathbf{L}_{k}^{e^{-}}$ using $\widehat{\mathbf{F}}\cdot\mathbf{n}\psi_{k}$
15:   for each element $e$ do
16:    for each quadrature point $i$ do
@@
21:      Compute $\mathbf{F}:=[U(\eta+h),V(\eta+h)]^{T}$
22:      for each modal basis $k=1,2,3$ do
23:       Accumulate $\mathbf{L}^{e}_{k}$ with $R(e)\psi_{k}$ and $\nabla\psi_{k}\cdot\mathbf{F}$
24:   for each element $e$ do
25:    for each mode $k$ do
26:      $(\zeta^{e})^{s+1}:=(\zeta^{e})^{s}+\Delta t\;\mathbf{L}^{e}$
27:   $\eta^{s}:=\eta^{s+1}$
@@
1: procedure CG_Momentum_Timestep($s$)
2:   for each element $e$ do
3:    $n_{1}:=N(e,1)$
4:    $n_{2}:=N(e,2)$
5:    $n_{3}:=N(e,3)$
6:    for $i=1,2,3$ do
7:      $\eta^{s}_{i}:=\eta^{s}(\mathbf{x}_{n_{i}})$
8:      $\eta^{s+1}_{i}:=\eta^{s+1}(\mathbf{x}_{n_{i}})$
9:    for $i=1,2,3$ do
10:      $H^{s}_{i}:=\eta^{s}_{i}+h(\mathbf{x}_{n_{i}})$
11:      $H^{s+1}_{i}:=\eta^{s+1}_{i}+h(\mathbf{x}_{n_{i}})$
12:    Compute elemental $\alpha,\beta_{\text{barotropic}}$ from $\eta^{s}_{i},\eta^{s+1}_{i},P^{s}(\mathbf{x}_{i}),P^{s+1}(\mathbf{x}_{i}),\;i=1,2,3$
13:    Compute elemental $\alpha,\beta_{\text{advect}}$ from $u_{x}^{n},u_{y}^{n}$
14:    for $i=1,2,3$ do $\triangleright$ Global assembly
15:      $\mathbf{X}_{n_{i}}:=\mathbf{X}_{n_{i}}+\alpha_{\text{barotropic}}+\alpha_{\text{advect}}$
16:      $\mathbf{Y}_{n_{i}}:=\mathbf{Y}_{n_{i}}+\beta_{\text{barotropic}}+\beta_{\text{advect}}$
                              $\triangleright$  Assembly of the load vectors
17:   for each node $i$ do
18:    Compute nodal $\alpha,\beta_{\text{fric}}$ using $H^{s}_{i},H^{s+1}_{i},u_{x}^{s},u_{x}^{s+1},u_{y}^{s},u_{y}^{s+1}$
19:    Compute nodal $\alpha,\beta_{\text{coriolis}}$ using $u_{x}^{s},u_{x}^{s+1},u_{y}^{s},u_{y}^{s+1}$
20:    $\mathbf{X}_{i}:=\mathbf{X}_{i}+\alpha_{\text{fric}}+\alpha_{\text{coriolis}}$
21:    $\mathbf{Y}_{i}:=\mathbf{Y}_{i}+\beta_{\text{fric}}+\beta_{\text{coriolis}}$
22:   for each node $i$ do
23:    Compute $\gamma_{\text{fric}},\gamma_{\text{coriolis}}$
24:    Solve the system for $u_{x}^{s+1}(\mathbf{x}_{i}),u_{y}^{s+1}(\mathbf{x}_{i})$
@@ 4.2 DG approach
Wetting and drying in DG-SWEM is considerably simpler as we only have to deal with elements and not nodes [bunya2009wetting]. It is constructed such that elemental mass is always conserved, and momentum whenever possible. For each element we apply the PositiveDepth() procedure which performs the following:
@@ 4.3 Combined approach
procedure PositiveDepth
   for each element $e$ do
@@
    $\overline{H}:=(H_{1}+H_{2}+H_{3})/3$
    if $H_{i}>H_{0},\;\forall i=1,2,3$ then
      $\omega_{e}:=1$
      for $i=1,2,3$ do
       $\Theta_{i}:=1$
      continue
    else if $\overline{H}<H_{0}$ then
      for $i=1,2,3$ do
       $\zeta_{i}:=\overline{H}-h_{N_{e,i}}$
      $\omega_{e}:=0$
    else
      Redistribute mass among $\zeta_{i}$ such that $\zeta_{i}+h_{(N(e,i)}\geqslant H_{0}$
      for $i=1,2,3$ do
@@
       $(u_{y})_{N_{e,i}}:=0$
    Reproject nodal $\zeta_{1},\zeta_{2},\zeta_{3}$ onto $\bm{\zeta}^{e}$
@@ 5 Incorporating Rainfall Data
During tropical cyclones, rainfall patterns in the vicinity of the storm exhibit defined structures that can be exploited by so-called parametric rainfall models. The basic idea behind this approach is to construct rainfall fields using simple analytic expressions based on a small number of (predicted or observed) storm parameters that are made publicly available by the National Hurricane Center (NHC) pre and post storm events. This type of approach has been implemented in the code based on the so-called R-CLIPER (Rainfall CLImatology and PERsistence) and the Interagency Performance Evaluation Task Force Rainfall Analysis (IPET) model, which computes rainfall rates $R(r)$ at any point in the grid, where $r$ is the distance from that point to the storm center. As an example, the R-CLIPER datapath is illustrated in Figure 3; see [Tuleya:2007, Brackins2020-xp] for more details on general parametric schemes. Note that the primary inputs of the model are the reported storm centers and maximum wind velocities, which are generally provided at 6-hour intervals by the NHC. Within DG-SWEM, linear interpolation is used to obtain the storm parameters at the model time step, and the rainfall rate at each finite element mesh node is computed based on its radial distance from the storm center.
@@
Alternatively, the program should also accept observed or reanalysis rainfall data in hindcasting scenarios in the GRIB2 format [grib2] which is read during each timestep. As this type of data is given with spatially varying distributions over the ocean and land, this data is readily interpolated onto finite element meshes. This option has been implemented in DG-SWEM and will be incorporated into ADCIRC in the future.
6 Numerical Experiments
@@
In this section, we thoroughly evaluate the developed solver by performing extensive numerical experiments that are designed to validate the solver. First, we consider simplified test cases where analytic solutions exist in simple rectangular and annular geometries. This is followed by the classical Shinnecock Inlet test case, a modified lake-at-rest [leveque1998balancing] test case, and a relatively new compound flood modeling benchmark of the Neches River. Last, we perform large-scale validation test cases for Hurricanes Ike (2008) and Harvey (2017) where we simulate the entire western North Atlantic Ocean, Caribbean Sea, and Gulf of Mexico.
@@ 6.1 Lynch and Gray test case
To verify the numerical accuracy of the scheme, we investigate the convergence of the error as we refine the mesh spacing $h$. This test case was first described in [Lynch1978-ih] which also presented an analytical steady-state solution to be used in the verification. This is for the linearized version of the SWE, obtained by neglecting the advection terms and linearizing the friction term:
@@ 6.5 Neches River
Figure 12: Computational domain of the Neches river case. Observation gauges are marked as red circles and boundaries are highlighted in yellow [loveland2021developing].
@@
The next test case is based on an approach to model compound flooding through river-coastal ocean interaction. In this case, we consider the Lower Neches River watershed near the Texas-Louisiana border as shown in Figure 12. During Hurricane Harvey (2017), this watershed was subjected to extreme rainfall and large portions of its surrounding areas were inundated during and after the event, see e.g., [watson2018characterization] for details on the flooding during this event. The model domain has been extracted from the ADCIRC mesh used in the Hurricane Harvey case in Section 6.6 [hope2013hindcast], and contains 122,839 elements and 62,075 nodes and is shown in Figure 12. Hence, the highly detailed unstructured mesh, bathymetry and topography, as well as parameters such as Manning’s $n$ are preserved. This event was also studied using ADCIRC and compared to HEC-RAS in [loveland2021developing] and the present study uses the same input data used in Loveland et al.. We only present key features of this model and refer readers to the original publication and [wichitrnithed2024discontinuous] where we extended the DG-SWEM model for greater detail.
As in [loveland2021developing], river flow data was extracted from a validated HEC-RAS model and a USGS gauge and implemented as flux boundary condition at the upstream. At the downstream end of the river which terminates into Sabine Lake, a time-varying elevation boundary condition was applied based on the same HEC-RAS model and the closest NOAA gauge. The locations of these inputs and the gauges used for validation are shown in Figure 12.
Comparisons of outputs from CG, DG-CG, as well as observed data from August 20 to September 1, 2017 are shown in Figure 13. We observe underprediction from the CG solver in all cases except at the Rainbow Bridge. We also observe that in those cases, the output from DG-CG is higher and closer to peak observations. At Rainbow Bridge, both solvers match closely (while slightly overpredicting) up to the end of observed data. That the DG scheme is able to better handle flows with high advection than ADCIRC has also been demonstrated in other studies [dawson2011discontinuous, wichitrnithed2024discontinuous]. The mismatch between wet/dry elevation values at the beginning of each solver can be attributed to differences in wetting-and-drying criteria as described in Section 4.
@@ 6.6 Hurricane Ike (2008) and Hurricane Harvey (2017)
In Figure 14, the computational mesh used is shown. This mesh consists of 2 million nodes and contains a recently updated version of the detailed floodplain, with a smallest element size of 120 meters [Contreras2023-qi]. Manning’s $n$ friction parameter is variable throughout the domain and was determined from observed datas [Contreras2023-qi]. As shown in Figure 14, the mesh is highly unstructured with increased refinement in the Texas-Louisiana coastal region. For Hurricane Ike, we use a proprietary wind data in the OWI format from OceanWeather Inc., see Section 4.1 in [hope2013hindcast] for further details on the wind data. For Hurricane Harvey, we use a more refined version of the mesh which contains around 8 million nodes; this is needed to cover the floodplain region where Harvey primarily affected. Wind data in this case is obtained from the NHC Hurricane Database (HURDAT) in the Best Track format. Lastly, for both hurricanes, we obtain tides using OceanMesh2D [Pringle2018], which in turn utilizes the TPXO9 tidal model [Egbert2002] to define a periodic forcing based on the Q1, O1, P1, K1, N2, M2, S2, and K2 tidal constituents. Additionally, we apply a elevation boundary condition corresponding to these tides along the 60 meridian, i.e., the eastern portion of the mesh shown in Figure 14.
@@
Water elevation comparison for the Harvey case is shown in in Figure 16. We can clearly observe the effect of rain in the DG-CG case which raises the water level following the peak surge. This is most apparent at Aransas Wildlife Refuge (8774230) and Sargent (8772985). In the former, the secondary peak is delayed in DG-CG but matches the actual peak level better than CG without rain. In the latter, both solvers underpredict initial water levels, but DG-CG eventually matches the peak level. On the other hand, the parametric rain approach can also lead to overprediction as seen in the Packery Channel station. We also compare the simulated maximum water levels with the observed high water marks (HWM) [hwm] located in the region around Harvey’s landfall (Figure 17). Note that data counts differ between CG and DG-CG because we only count nodes that are wet. We see a marked improvement in the DG-CG case, indicating the impact of precipitation in compound flooding. A limitation in DG-CG wetting and drying for this case is that it was necessary to turn off momentum advection terms to prevent instability. Although these terms do not make a significant impact in this case, we do aim to make the DG-CG wetting and drying scheme more robust so as to handle highly refined meshes such as this one.
@@ 7.1.1 Optimization
As mentioned, the higher degrees of freedom in the DG method unavoidably requires more time to compute than the CG scheme. On the other hand, the higher degree of data parallelism from having most computations be element-local makes it attractive to use vector instructions. The general idea and motivation is outlined in [Brus2017-ve] which involves significant loop reordering and blocking. To keep the code maintainable as part of ADCIRC, we opt to use a simpler approach and try to rely on automatic vectorization and basic optimizations as much as possible [Unknown2010-wl]. Wherever possible, loops with independent iterations and large trip count (e.g. number of elements) are prepended with the OpenMP !\$omp simd pragma, and inner loops are unrolled. The compiled code and assembly is checked using Intel Advisor to verify if the vectorization took place. Vectorizing the slope limiter requires a little more effort, as the automatic vectorizations were only partial. We manually performed explicit blocking on the number of elements and created vector versions of the local arrays private to each element. Without significant restructuring, the interior edge integration routine is not very suitable to vectorization since multiple edges can modify the same element. Nevertheless, we still observe significant improvement from inlining the numerical flux function by moving it into the same module.